OpenStudy (anonymous):

Eva created a coffee blend for her restaurant by mixing Kenya and French Roast coffee beans. The Kenya beans cost $15 per pound and the French Roast costs $8 per pound. She bought a total of 10 pounds of coffee for a total of $97.50. How much of each type of bean did she buy? Select one: a. 2.5 lbs Kenya and 7.5 lbs French Roast b. 3 lbs Kenya and 7 lbs French Roast c. 1.5 lbs Kenya and 8.5 lbs French Roast d. 4 lbs Kenya and 6 lbs French Roast

OpenStudy (anonymous):

B and D are wrong.

OpenStudy (anonymous):

Answer is A

OpenStudy (anonymous):

2.5 x 15 = 37.5 7.5 x 8 = 60 60 + 37.5 = 97.5

OpenStudy (yanasidlinskiy):

x=kenya beans pounds y=french roasts pounds Now, create a system of equations. You have 2: 1. x+y=10 2. 15x+8y=97.50 Then, solve system of equations. y=10-x 15x+8(10-x)=97.50 7x+80=97.50 7x=17.5 x=2.5. plug into y=10-x. y=7.5. 2.5 lbs Kenya and 7.5 lbs French roast.

OpenStudy (anonymous):

@Sarah.Counters Understand?

OpenStudy (anonymous):

yeah and thanks

OpenStudy (yanasidlinskiy):

:D

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

Got more questions?

OpenStudy (anonymous):

Or was that your only one?

OpenStudy (anonymous):

got more